OpenStudy (anonymous):

Find all solutions of the equation 3csc^2x - 4 = 0

OpenStudy (mimi_x3):

\[3csc^{2}x = 4\] \[ csc^2(x) = \frac{4}{3}\] \[\frac{1}{sin^{2}x} = \frac{4}{3}\] \[ 3 = 4sin^{2}x => 4sin^{2}x = 3 => sin^{2}x = \frac{3}{4}\]

OpenStudy (mimi_x3):

You able to do it now?

OpenStudy (anonymous):

Um. So would all of the solutions be: 60 + 360n ?

terenzreignz (terenzreignz):

Sure... are you sure it isn't +180n instead?

OpenStudy (anonymous):

Wait why would it be +180n?

terenzreignz (terenzreignz):

Because..., for instance, 60 + 180 = 240 Try sin(240)

terenzreignz (terenzreignz):

Right, I know it's going to give you something negative, but remember, we're looking for the solutions to \[\large \sin^2(x)=\frac{3}{4}\] and not just \[\large \sin(x) = \sqrt{\frac34} =\frac{\sqrt3}{2}\]

terenzreignz (terenzreignz):

So actually, a more specific way to write it would be finding the solutions to \[\Large \sin(x) = \pm \frac{\sqrt3}{2}\]

OpenStudy (anonymous):

would the answer for negative \[\sqrt{3}/2\] be the same as its positive?

terenzreignz (terenzreignz):

No... but we're looking for the solutions to the equation \[\large \sin^2(x) = \frac34\] right? Which means that if we take the square root of both sides... \[\Large \sin(x) = \pm\sqrt{\frac34}\]we have to consider both the positive and negative square roots, since we are going to square it...

terenzreignz (terenzreignz):

you can also look at it this way... \[\Large \sin^2(x) - \frac34 = 0\\\\\\\\\Large let \ u=\sin(x) \] \[\Large u^2 - \frac34 = 0\] \[\Large \left(u-\sqrt \frac34\right)\left(u+\sqrt \frac34\right)=0\]

OpenStudy (anonymous):

Oh ok. I still don't really get this concept though. For example, if I have.... 4 cos^2 x - 1 = 0 How would I find all of the solutions?

terenzreignz (terenzreignz):

I'm sure it would be a very similar problem... if it makes things easier, you can always let \(\Large u = \cos(x)\) So it becomes \[\Large 4u^2 -1 = 0\] and then solve for u first...

OpenStudy (anonymous):

What would you do after you solve for u and get +/- 1/2 ? How would you go about in finding all of the solutions?

terenzreignz (terenzreignz):

Well, replace u again with cos(x) \[\huge \cos(x) = \pm \frac12\] and solve for all values of x whose cosines are either 1/2 or -1/2

OpenStudy (anonymous):

Would that be 60 + 300n?

terenzreignz (terenzreignz):

300n? You mean 360n?

OpenStudy (anonymous):

Wait why would it be 360n?

terenzreignz (terenzreignz):

Why would it be 300n? 360n is more understandable, as rotating 360 degrees essentially brings it back to the same angle... 300n? at n = 1, it already doesn't hold... cos(60 + 300) = cos(360) = 1

terenzreignz (terenzreignz):

But actually, much like earlier, it's 60 + 180n

terenzreignz (terenzreignz):

and also 120 + 180n

OpenStudy (anonymous):

Oh ok. I see. Thank you!!!!

terenzreignz (terenzreignz):

No problem :)